OpenStudy (lsdxmdmaddicthc):

A regular hexagon has a perimeter of 60 in. What is the hexagon's area?

OpenStudy (pinklion23):

easy i will help

OpenStudy (lsdxmdmaddicthc):

@sleepyjess

OpenStudy (lsdxmdmaddicthc):

I want to learn how to do it with my brain.

OpenStudy (pinklion23):

medal pls?

OpenStudy (anonymous):

Ok

OpenStudy (sleepyjess):

Do you know the formula for area of a hexagon?

OpenStudy (lsdxmdmaddicthc):

No I don't, sleepyjess

OpenStudy (sleepyjess):

Okay, the formula is \(\dfrac{3\sqrt 3}{2}*a^2\). In this formula, a is the length of one side

OpenStudy (lsdxmdmaddicthc):

Area = (3*3(sqrt)/2) * a(sq)

OpenStudy (lsdxmdmaddicthc):

60/6 = 10 units each side

OpenStudy (lsdxmdmaddicthc):

inches*

OpenStudy (sleepyjess):

Yep, so \(\dfrac{3\sqrt 3}2 *1-^2\)

OpenStudy (sleepyjess):

That should be \(10^2\)

OpenStudy (lsdxmdmaddicthc):

(3*(sqrt)3/2)*10(sq) (3*(sqrt)3/2)*100 (5.2/2)*100

OpenStudy (lsdxmdmaddicthc):

2.6*100 = 26

OpenStudy (sleepyjess):

Does it tell you to round to a certain decimal place?

OpenStudy (lsdxmdmaddicthc):

No I just rounded when doing the math, I rounded to the nearest hundreths

OpenStudy (sleepyjess):

Okay, normally I keep it at the nearest hundredths, but you actually rounded to the nearest tenth

OpenStudy (lsdxmdmaddicthc):

It's multiple choice here's the answers : 75 (sqrt of 3) 150 (sqrt of 3) 300 (sqrt of 3) 600 (sqrt of 3)

OpenStudy (sleepyjess):

Oh, okay, they don't want you to multiply the \(\sqrt 3\). So we multiply the 3 *100, that will be \(\dfrac{300\sqrt 3}{2}\) Then just divide the 300 by 2

OpenStudy (lsdxmdmaddicthc):

Okay So 150 (sqrt 3)

OpenStudy (lsdxmdmaddicthc):

Thanks again sleepyjess! You're a great help I'm a fan

OpenStudy (sleepyjess):

No problem! :)

OpenStudy (sleepyjess):

Thank you for participating instead of just using an online calculator :)

OpenStudy (lsdxmdmaddicthc):

Haha, well I actually enjoy mathematics it's a puzzle. I just needed a guide.

OpenStudy (lsdxmdmaddicthc):

Haha, well I actually enjoy mathematics it's a puzzle. I just needed a guide.